OpenStudy (anonymous):

Calculate the range of a projectile using the initial speed of 18 m/s and angle of 75 degrees. Use g=9.8 m/s for the calculation... Please help..

OpenStudy (anonymous):

Formula for Range of Projectile: R=vo² sin2 Θ /g vo=18m/s 2Θ=150 g=9.8m/s^2 R=(18^2)*sin150/9.8 R=324*0.5/9.8 R=162/9.8 R=16.5 meters